(16 points) Find y as a function of x if y + 25y = 0, y(0) = -7, y (O) = -15, y (0) = 100. y(x) =
Answer #1

Given differential equation is y+25y= 0 d (D* +25D) y = 0 where D= Auxiliary equation is m +25m= 0 m(m2 +25) = 0 m= 0, m =